# Forget BIT, It is All about TOKEN: Towards Semantic Information Theory for LLMs

### v1

@@ V-E Memory Capacity of the Transformer
The statistical physics approaches, such as spin glass model and replica method, have been widely used to analyze the performance of signal processing, coding, and satisfiability (SAT) problems [91]. In a series of landmark papers [53, 54, 55], Gardner investigated the memory capacity of the classical Hopfield network [92] by applying the replica method, which is referred to as Gardner capacity afterwards.
@@ V-F Semantic Information Theoretic Measure for the Transformer
The directed information $I(S_{1:n}\to U_{n+1:t};\Phi)$ can be represented by KL divergence as follows
$I(S_{1:n}\to U_{n+1:t};\Phi)=D_{KL}\left(P(S_{1:n},U_{n+1,t})\|P(S_{1:n})\prod_{j=n+1}^{t}P(U_{j}|U_{n+1:j})\right).$ (77)
@@ V-G1 Mamba/Mamba2
$\left\{\begin{aligned} \mathbf{u}_{t}&=\mathbf{A}_{t}\mathbf{u}_{t-1}+\mathbf{B}_{t}\mathbf{s}_{t};\\ \mathbf{y}_{t}&=\mathbf{C}\mathbf{u}_{t}.\end{aligned}\right$. (78)
Clearly, the SSM is a special case of the AR-LLM in Definition 13, which exactly belongs to linear TV-VAR models [50]. The linear TV-VAR model is widely used in time series analysis for economics and finance [97, 98]. Therefore, the developed parameter estimation method may be applicable to improve the performance of Mamba/Mamba2. Because there lacks the bilinear model of semantic relevance, it is not difficult to understand that the performance of Mamba/Mamba2 could be worse than Transformer. However, the Mamba/Mamba2 architectures inspire us to consider other forms of AR-LLM which may have a similar performance as Transformer but much lower computation complexity. Based on the improved Mamba2 [99], Qwen3-Next is the first LLM which implements the hybrid attention scheme. [^4] https://qwen.ai/blog?from=research.latest-advancements-list&id=4074cca80393150c248e508aa62983f9cb7d27cd& The Transformer, however, is different from linear TV-VAR model because $\pi_{tj}$ introduces a non-linear relation, i.e., the $\mathrm{softmax}$ function over a bilinear form of $\mathbf{u}_{t}$ and $\mathbf{u}_{j}$.
@@ V-G2 LLaDA
$\mathcal{L}(\Phi)=-\mathbb{E}_{\tau,U_{1:T}^{\tau},U_{1:T}^{0}}\left\{\frac{1}{\tau}\sum_{t=1}^{T}\mathbf{1}(U_{t}^{\tau}=M)\log P(U_{t}^{0}|U_{1:T}^{\tau};\Phi)\right\},$ (79)
where $M$ denote the masked token. The transformer without causal mask is used as the core component to predict the masked tokens. Evidently, while LLaDA is fundamentally built upon a diffusion framework, the AR-LLM remains central to the task of masked token prediction in LLaDA.
@@ VI Conclusions
Drawing from the theory of rate-distortion function, directed information, and Granger causality, this paper aims to uncover the semantic information-theoretic principles underlying LLMs. We discussed the structure-agnostic information-theoretic measures, the token-level semantic embedding, and the general definition of AR-LLM, from which the Transformer architecture and its performance have been derived theoretically. Our theory indicates that the capabilities of current LLMs remain within the scope of Granger causality. How to achieve the counterfactual reasoning and system 2 reasoning abilities [100, 101], remains a formidable challenge. Consequently, our semantic information theory framework provides a lens through which many experimentally observations can be explained, which also paves the way for unlocking the full potential of LLMs.
@@ Acknowledgment
I am grateful to T. Wu, X. Niu, K. Zhang, C. Zhang, Y. Lan, Z. Zhong, B. Chen, and Q. Zhang for productive discussions.

### v2

@@ V-E Memory Capacity of the Transformer
The statistical physics approaches, such as spin glass model and replica method, have been widely used to analyze the performance of signal processing, coding, and satisfiability (SAT) problems [95, 96]. In a series of landmark papers [53, 54, 55], Gardner investigated the memory capacity of the classical Hopfield network [97] by applying the replica method, which is referred to as Gardner capacity afterwards.
@@ V-F Semantic Information Theoretic Measure for the Transformer
The directed information $I(S_{1:n}\to U_{n+1:i};\Phi)$ can be represented by KL divergence as follows
$I(S_{1:n}\to U_{n+1:i};\Phi)=D_{KL}\left(P(S_{1:n},U_{n+1,i})\|P(S_{1:n})\prod_{j=n+1}^{i}P(U_{j}|U_{n+1:j})\right).$ (80)
@@ V-G1 Mamba/Mamba2
$\left\{\begin{aligned} \mathbf{u}_{i}&=\mathbf{A}_{i}\mathbf{u}_{i-1}+\mathbf{B}_{i}\mathbf{s}_{i};\\ \mathbf{y}_{i}&=\mathbf{C}\mathbf{u}_{i}.\end{aligned}\right$. (81)
Clearly, the SSM is a special case of the AR-LLM in Definition 13, which exactly belongs to linear TV-VAR models [50]. The linear TV-VAR model is widely used in time series analysis for economics and finance [102, 103]. Therefore, the developed parameter estimation method may be applicable to improve the performance of Mamba/Mamba2. Because there lacks the bilinear model of semantic relevance, it is not difficult to understand that the performance of Mamba/Mamba2 could be worse than Transformer. However, the Mamba/Mamba2 architectures inspire us to consider other forms of AR-LLM which may have a similar performance as Transformer but much lower computation complexity. Based on the improved Mamba2 [104], Qwen3-Next is the first LLM which implements the hybrid attention scheme. [^4] https://qwen.ai/blog?from=research.latest-advancements-list&id=4074cca80393150c248e508aa62983f9cb7d27cd& The Transformer, however, is different from linear TV-VAR model because $\pi_{ij}$ introduces a non-linear relation, i.e., the $\mathrm{softmax}$ function over a bilinear form of $\mathbf{u}_{i}$ and $\mathbf{u}_{j}$.
@@ V-G2 LLaDA
$\mathcal{L}(\Phi)=-\mathbb{E}_{\tau,U_{1:N}^{\tau},U_{1:N}^{0}}\left\{\frac{1}{\tau}\sum_{i=1}^{N}\mathbf{1}(U_{i}^{\tau}=\mathsf{M})\log P(U_{i}^{0}|U_{1:N}^{\tau};\Phi)\right\},$ (82)
where $\mathsf{M}$ denote the masked token. The transformer without causal mask is used as the core component to predict the masked tokens. Evidently, while LLaDA is fundamentally built upon a diffusion framework, the AR-LLM remains central to the task of masked token prediction in LLaDA.
@@ VI Conclusions
Drawing from the theory of rate-distortion function, directed information, and Granger causality, this paper aims to uncover the semantic information-theoretic principles underlying LLMs. We discussed the architecture-agnostic information-theoretic measures, the token-level semantic embedding, and the general definition of AR-LLM, from which the Transformer architecture and its performance have been derived theoretically. Our theory indicates that the capabilities of current LLMs remain within the scope of Granger causality. How to achieve the counterfactual reasoning and system 2 reasoning abilities [105, 106], remains a formidable challenge. Consequently, our semantic information theory framework provides a lens through which many experimentally observations can be explained, which also paves the way for unlocking the full potential of LLMs.
@@ Acknowledgment
I am grateful to T. Wu, X. Niu, K. Zhang, L. Li, C. Zhang, Y. Lan, G. Zhang, Z. Zhong, B. Chen, and Q. Zhang for productive discussions.